The flower garden has the shape of a right triangle. 17ft of a perennial border forms the hypotenuse of the triangle, and one leg is 7ft longer than the other leg. Find the lengths of the legs.
Answer by cleomenius(959)

Since this is a right triangle, we can use a^2 + b^2 = c^2
Let one leg be x.
Let the other leg be x + 7.
x^2 + (x + 7)^2 = 289
x^2 + x^2 + 14x + 49 = 289
2x^2 + 14x - 240 = 0
2 ( x^2 + 7x - 120) = 0
2 (x - 8) ( x + 15) = 0
x will = 8, one leg of the triangle.
8 + 7 = 15, the other leg of the triangle.
8^2 + 15^2 = 17^2
64 + 225 = 289, The results do check.
Cleomenius.